OpenStudy (czarluc):

divide 30 into two parts so that the sum of their squares is a minimum

OpenStudy (solomonzelman):

What course is this?

OpenStudy (czarluc):

uhm quadratic equations

OpenStudy (p0sitr0n):

a+b=30 a^2+b^2=p(a,b) minimize p basically

OpenStudy (solomonzelman):

well, just logically you know that for example: \(\large\color{red}{ 40^2+1^2>20^2+20^2 }\)

OpenStudy (czarluc):

ok ok

OpenStudy (solomonzelman):

this is an optimization problem? because accroding to what Positron posted it really lokes like one.

OpenStudy (czarluc):

it was asked to be solved using quadratic equations

OpenStudy (solomonzelman):

I lost connection. Love this place. Anyway... I really don't know how to actually solve (without using a pure logic) if I am not taking the derivative of your function when you solve the a^2+b^2=S for one variable. Like I was thinking of basic calc problem, but not just using a quadratic equation.

OpenStudy (czarluc):

haha yeah, I often loose connections too. they say it would be like x^2+(30-x)^2=0 but I don't understand much of it

OpenStudy (solomonzelman):

and how do they know to set it to zero. i see they are plugging b-30 for a, and re-writing the Bs with Xs.

OpenStudy (solomonzelman):

I don't understand it either. I think it is not the correct way....

OpenStudy (czarluc):

Oh, and I remeber that it should be used with the extreme values lik V(h,k)

OpenStudy (xapproachesinfinity):

this is calculus problem

OpenStudy (solomonzelman):

:)

OpenStudy (zarkon):

you don't need calculus to do this problem

OpenStudy (xapproachesinfinity):

well consider the two pat are x and y since x+y=30====> y=x-30 then S=x^2+(x-30)^2

OpenStudy (xapproachesinfinity):

that may be but the problem is asking for calculus here i would guess

OpenStudy (solomonzelman):

or just logic?

OpenStudy (solomonzelman):

Again, it is logical, and makes sens that: \(\large\color{black}{ 25^2+25^2<50^2+1^2~.}\)

OpenStudy (zarkon):

\[x^2+(30-x)^2=2x^2-60x+900\] this is a quadratic that goes up. it has a min at the vertex \[\frac{-(-60)}{2\cdot 2}=\frac{60}{4}=15\]

OpenStudy (xapproachesinfinity):

hmm yes! quadratic solves the problem too

OpenStudy (solomonzelman):

ty @Zarkon (sorry for tagging)

OpenStudy (czarluc):

Oh ok I get it! thanks everyone!